OpenStudy (anonymous):

Which of the following are geometric sequences?Check all that apply. A. 2, -2, 2, -2, 2, -2, 2 B. 1, 2, 4, 8, 16, 32 C. 1, 4, 9, 16, 25, 36, 49 D. 10, 5, 2.5, 1.25, 0.625, 0.3125

OpenStudy (anonymous):

first one is \[2\times (-1)^{n+1}\] go geometric. notice that each ratio is -1

OpenStudy (anonymous):

second one is \[2^{n-1}\] also geometric. notice again that common ratio is 2

OpenStudy (anonymous):

third one is \[n^2\] not geometric because for example \[\frac{25}{16}\neq \frac{36}{25}\]

OpenStudy (anonymous):

last one ... take the ratio of any two successive terms. if those ratios are the same, then it is geometric.